OpenStudy (anonymous):

An electronics store makes a profit of $72 for every television sold and $90 for every computer sold. The manager’s target is to make at least $360 a day on sales from televisions and computers. Write a linear inequality and graph the solutions. What are three possible solutions to the problem?

OpenStudy (texaschic101):

72t + 90c >= 360 Three possible solutions can be found by subbing in any number for t and solving for c (1) 72 (2) + 90c > = 360 144 + 90c >= 360 90c >= 360 - 144 90c >= 216 c >= 216/90 c >= 2.4 one solution is : t = 2 and c >= 2.4 (2) 72 (3) + 90c >= 360 216 + 90c >= 360 90c >= 360 - 216 90c >= 144 c >= 144/90 c >= 1.6 solution : t = 3 and c >= 1.6 (3) 72(1) + 90c >= 360 72 + 90c >= 360 90c >= 360 - 72 90c >= 288 c >= 288/90 c >= 3.2 solution : t = 1, and c >= 3.2

OpenStudy (anonymous):

in my choices none of them have a decimal.

OpenStudy (texaschic101):

what are the choices ?

OpenStudy (anonymous):

A.)(5,2) (3,3) (1,4) B.)(4,0) (2,2) (1,1) C.)(3,1) (3,3) (1,4) D.)(4,0) (3,3) (1,4)

OpenStudy (texaschic101):

I believe your answer is A